# BIMBA: Selective-Scan Compression for Long-Range Video Question Answering

### v1

@@ Abstract
Video Question Answering (VQA) in long videos poses the key challenge of extracting relevant information and modeling long-range dependencies from many redundant frames. The self-attention mechanism provides a general solution for sequence modeling, but it has a prohibitive cost when applied to a massive number of spatiotemporal tokens in long videos. Most prior methods rely on compression strategies to lower the computational cost, such as reducing the input length via sparse frame sampling or compressing the output sequence passed to the large language model (LLM) via space-time pooling. However, these naive approaches over-represent redundant information and often miss salient events or fast-occurring space-time patterns. In this work, we introduce BIMBA, an efficient state-space model to handle long-form videos. Our model leverages the selective scan algorithm to learn to effectively select critical information from high-dimensional video and transform it into a reduced token sequence for efficient LLM processing. Extensive experiments demonstrate that BIMBA  achieves state-of-the-art accuracy on multiple long-form VQA benchmarks, including PerceptionTest, NExT-QA, EgoSchema, VNBench, LongVideoBench, and Video-MME. Code, and models are publicly available at https://sites.google.com/view/bimba-mllm.
@@ 1 Introduction
However, most video MLLMs have been applied to short video inputs spanning a few seconds. In this case, a simple but effective strategy involves extracting “tokens” from the individual frames of the video using a pretrained image encoder to produce outputs aligned to the space of an LLM [53, 46, 36, 35]. The concatenation of the tokens extracted from the frames is then passed to the LLM for semantic processing and language generation [3, 36, 66]. This strategy breaks down when the video is very long, e.g., when it spans minutes or hours, because of the quadratic cost of the self-attention operation needed to process the extracted visual tokens. In addition, naive concatenation of frame-level tokens for a long video would flood the LLM with an exceedingly long and redundant sequence of features. For example, the LLAMA-3.2 [53] image encoder outputs 1600 $\sim$ 6400 tokens per image. Therefore, encoding just 128 frames would produce a sequence of 205K $\sim$ 820K tokens, which is impractical to process even by modern GPUs.
@@
Although SSM and Mamba models have been used before for image analysis [90, 57, 57, 8], image-language modeling [59, 86, 26, 34] and video understanding [77, 41, 50, 6, 27, 28], our proposed architecture differs from these prior works by introducing a few simple but effective contributions. First, we introduce an SSM-based compression module called a spatiotemporal token selector, which can take a large number of spatiotemporal video tokens as input and output a significantly smaller number of compressed tokens only containing important information by utilizing the selective-scan layers inside the module. We achieve this by introducing a smaller number of visual queries and jointly modeling them with the spatiotemporal video tokens. Second, we propose an improved method for concatenating visual and spatiotemporal tokens by interleaving them at equal intervals. The standard approach of appending tokens at the end can introduce positional bias, where end-sequence tokens disproportionately influence query tokens. Our interleaved positioning mitigates this bias, influencing queries more uniformly across the entire video sequence. Finally, we use a bidirectional selective-scan mechanism [41] that is more effective in capturing 2D/3D spatiotemporal structures in video compared to the standard selective scan, which is better suited for 1D sequence modeling.
We experiment on several long video question answering datasets including PerceptionTest [56], NExT-QA [73], EgoSchema [52], VNBench [87], LongVideoBench [72], Video-MME [16], and show state-of-the-art results on all of these datasets.
@@ 3 Background: Mamba Model
The S4 model is a Linear Time-Invariant (LTI) system with parameters that remain constant and are independent of the input. This can be suboptimal, especially for modeling long-term dependencies in a sequence containing redundant and unimportant information. Recent work [18] proposes Selective Scan State Space (S6) to develop Mamba LLM [18]. In S6, the matrices $\mathbf{\overline{B}}$, $\mathbf{C}$, and $\mathbf{\Delta}$ are made input dependent and derived from input $x$ using linear layers as follows.
@@ 5.2 Evaluation Benchmarks
We evaluate our model on six diverse video question-answering benchmarks: Perception Test [56], NExT-QA [74], EgoSchema [52], VNBench [87], LongVideoBench [72], and Video-MME [16].
@@ 6.4 Comparison with the State of the Art
Model LLM Frames PerceptionTest NExT-QA EgoSchema VNBench LongVideoBench Video-MME
Duration 23 sec 44 sec 180 sec 180 sec 8 sec - 60 min 1 - 60 min
@@
LLoVi [81] GPT3.5 - - 66.30 52.00 - - -
VideoAgent [68] GPT4 - - 71.30 54.10 - - -
VideoTree [70] GPT4 - - 75.60 61.10 - - -
@@
LLaVA-NeXT (Video) [45] Vicuna-7B 64 46.13 67.66 41.66 - - 42.21
PLLaVA [76] Vicuna-7B 64 48.55 67.56 43.36 - - 42.13
BIMBA-LLaVA Vicuna-7B 64 52.61 72.35 52.31 - - 45.66
LLaMA-3.2 (Video) [53] LLaMA3.2-8B 64 53.16 73.72 55.21 - - 47.31
PLLaVA (LLaMA-3.2) [53] LLaMA3.2-8B 64 53.13 72.77 52.33 - - 46.56
BIMBA-LLaMA LLaMA3.2-8B 64 56.50 76.88 60.25 - - 50.11
@@
LLaMA-VID [42] Vicuna-7B 1fps 44.60 - 38.50 10.80 - -
Video-LLaVA [43] Vicuna-7B 8 44.30 62.60 38.40 - - 40.40
LLaVA-NeXT-Video [84] Vicuna-7B 32 48.80 - 43.90 20.10 43.50 46.50
PLLaVA [76] Vicuna-7B 16 - 68.17 45.16 - 40.20 44.25
VideoChat2 [40] Vicuna-7B 16 47.30 - 54.40 12.40 36.00 47.90
LongVA [83] Qwen2-7B 128 - 68.30 - 41.50 - 54.30
Video-LLaMA2 [10] Qwen2-7B 32 51.40 - 51.70 24.90 - 47.90
LLaVA-OneVision [35] Qwen2-7B 32 57.10 79.40 60.10 51.80 56.5 58.20
Kangaroo [48] LLaMA2-8B 64 - - 62.70 - - 56.00
Video-XL [63] Qwen2-7B 2048 - - - 61.60 49.50 55.50
LongVU [62] Qwen2-7B 1fps - - 67.60 - - 60.60
Qwen2-VL [67] Qwen2-7B 2fps - - - - 55.60 63.30
LLaVA-Video [85] Qwen2-7B 64 67.90 83.20 57.30 70.77 58.20 63.30
BIMBA-LLaVA Qwen2-7B 128 68.51 83.73 71.14 77.88 59.46 64.67
@@
Table 3: We compare BIMBA with state-of-the-art video MLLMs across six diverse video question-answering benchmarks. BIMBA-LLaVA achieves the highest performance on all datasets when using the Qwen2-7B LLM backbone (third section). Since different MLLMs leverage varying LLM backbones and training data, we also conduct a fair comparison by evaluating our model against four baselines trained using the same 370K instruction-tuning dataset and using Vicuna-7B and LLaMA3.2-8B LLM decoders (second section). In this evaluation setting, we observe that our model consistently outperforms all baselines across all benchmarks.
@@
In this section, we compare our model against state-of-the-art video MLLMs on six different video benchmarks, summarizing the results in Table 3. Since prior models rely on varying backbones and training datasets, to ensure a fair comparison, we first fine-tune four competitive models, as well as the two default variants of our BIMBA model on our 370K instruction-tuning dataset. These include PLLaVA [76], LLaVA-NeXT (Video) [45], LLaMA-3.2 [53] (Video) and PLLaVA (LLaMA-3.2), a variant of PLLaVA using the stronger LLaMA-3.2 [53] image MLLM. The results in the second section of the table show that our model achieves better results than all these baselines. In particular, we note that when using the same Vicuna-7B [11] model, BIMBA-LLaVA outperforms both PLLaVA and LLaVA-NeXT (Video) on all benchmarks, achieving an average accuracy of 55.23% compared to 49.42% for LLaVA-NeXT and 50.40% for PLLaVA. Likewise, our BIMBA-LLaMA variant surpasses LLaMA-3.2 (Video) by 3.09% on average and PLLaVA (LLaMA-3.2) by 4.21%. These results highlight the effectiveness of our selective-scan-based token compression mechanism, which can be seamlessly integrated into existing MLLMs to enhance performance across a wide range of video understanding tasks.
@@
Furthermore, since recent MLLMs [35, 48, 63, 62, 67, 85] have been trained with significantly larger instruction-tuning datasets and stronger LLM backbones, we also trained a variant of BIMBA-LLaVA using approximately 1.6 million video instruction samples and the Qwen2-7B LLM backbone. The results in the third section of Table 3 show that this BIMBA-LLaVA variant achieves state-of-the-art performance on multiple challenging video question-answering benchmarks, including NExT-QA [74], EgoSchema [52], VNBench [87], LongVideoBench [72], and Video-MME [16]. On the EgoSchema benchmark, our model surpasses the previous best method, LongVU [62], by 3.54%, demonstrating its superior ability to comprehend egocentric videos and handle questions that require long context understanding. Similarly, on VNBench, which focuses on needle-in-the-haystack questions, our approach outperforms LLaVA-Video [85] by 7.11%, highlighting its strong capability to extract key information from very long videos. Additionally, BIMBA-LLaVA (Qwen2-7B) substantially outperforms models based on proprietary LMMs. Specifically, it surpasses VideoTree [70] (GPT-4) by 7.7% on NExT-QA and 10.04% on EgoSchema. Finally, on benchmarks requiring long video comprehension, such as LongVideoBench [72] and Video-MME [16], our model sets a new state-of-the-art, further demonstrating its effectiveness in processing and understanding hour-long videos.
@@ 7 Conclusion
[…]
@@
BIMBA-LLaVA (Qwen2-7B) is built on LLaVA-Video [85], leveraging SigLIP [80] as the vision encoder and Qwen2-7B [12] as the LLM. It processes 128 video frames at a resolution of $336\times 336$, dividing each frame into $14\times 14$ patches, producing $128\times 24\times 24$ spatiotemporal tokens. These are compressed to $128\times 12\times 12$ before being passed to the LLM. Similar to the previous variant, the vision encoder remains frozen, while the multimodal projector, spatiotemporal token selector, and LLM are trained using LoRA.
@@
005050505010010010010015015015015000101010102020202030303030SecondsAccuracy Improvement(a) LLaVA Backbone on NeXT-QA
0050505050100100100100150150150150005555101010101515151520202020SecondsAccuracy Improvement(b) LLaMA Backbone on NeXT-QA

### v2

@@ Abstract
Video Question Answering (VQA) in long videos poses the key challenge of extracting relevant information and modeling long-range dependencies from many redundant frames. The self-attention mechanism provides a general solution for sequence modeling, but it has a prohibitive cost when applied to a massive number of spatiotemporal tokens in long videos. Most prior methods rely on compression strategies to lower the computational cost, such as reducing the input length via sparse frame sampling or compressing the output sequence passed to the large language model (LLM) via space-time pooling. However, these naive approaches over-represent redundant information and often miss salient events or fast-occurring space-time patterns. In this work, we introduce BIMBA, an efficient state-space model to handle long-form videos. Our model leverages the selective scan algorithm to learn to effectively select critical information from high-dimensional video and transform it into a reduced token sequence for efficient LLM processing. Extensive experiments demonstrate that BIMBA  achieves state-of-the-art accuracy on multiple long-form VQA benchmarks, including PerceptionTest, NExT-QA, EgoSchema, VNBench, LongVideoBench, Video-MME, and MLVU. Code and models are available at https://sites.google.com/view/bimba-mllm.
@@ 1 Introduction
However, most video MLLMs have been applied to short video inputs spanning a few seconds. In this case, a simple but effective strategy involves extracting “tokens” from the individual frames of the video using a pretrained image encoder [52, 3, 45, 34, 65, 35]. The concatenation of the tokens extracted from the frames is then passed to the LLM for semantic processing and language generation. This strategy breaks down when the video is very long, e.g., when it spans minutes or hours, because of the quadratic cost of the self-attention operation needed to process the extracted visual tokens. In addition, naive concatenation of frame-level tokens for a long video would flood the LLM with an exceedingly long and redundant sequence of features. For example, the LLAMA-3.2 [52] image encoder outputs 1600 $\sim$ 6400 tokens per image. Therefore, encoding just 128 frames would produce a sequence of 205K $\sim$ 820K tokens, which is impractical to process even by modern GPUs.
@@
Although SSM and Mamba models have been used before for image analysis [89, 56, 56, 8], image-language modeling [58, 84, 25, 33] and video understanding [76, 40, 49, 6, 26, 27], our proposed architecture differs from these prior works by introducing a few simple but effective contributions. First, we introduce an SSM-based compression module called a spatiotemporal token selector, which can take a large number of spatiotemporal video tokens as input and output a significantly smaller number of compressed tokens only containing important information by utilizing the selective-scan mechanism [17]. We achieve this by introducing a smaller number of visual queries and jointly modeling them with the spatiotemporal video tokens. Second, we propose an improved method for concatenating visual and spatiotemporal tokens by interleaving them at equal intervals. The standard approach of appending tokens at the end can introduce positional bias, where end-sequence tokens disproportionately influence query tokens. Our interleaved positioning mitigates this bias, influencing queries more uniformly across the entire video sequence. Finally, we use a bidirectional selective-scan strategy [40] that is more effective in capturing 2D/3D spatiotemporal structures in video compared to the standard selective scan, which is better suited for 1D sequence modeling.
We experiment on several long video question answering datasets including PerceptionTest [55], NExT-QA [72], EgoSchema [51], VNBench [85], LongVideoBench [71], Video-MME [15], MLVU [86], and show state-of-the-art results on all of these datasets.
@@ 3 Background: Mamba Model
The S4 model is a Linear Time-Invariant (LTI) system with parameters that remain constant and are independent of the input. This can be suboptimal, especially for modeling long-term dependencies in a sequence containing redundant and unimportant information. Recent work [17] proposes Selective Scan State Space (S6) to develop Mamba LLM [17]. In S6, the matrices $\mathbf{\overline{B}}$, $\mathbf{C}$, and $\mathbf{\Delta}$ are made input dependent and derived from input $x$ using linear layers.
@@ 5.2 Evaluation Benchmarks
We evaluate our model on seven diverse video question-answering benchmarks: Perception Test [55], NExT-QA [73], EgoSchema [51], VNBench [85], LongVideoBench [71], Video-MME [15], and MLVU [86].
@@ 6.4 Comparison with the State of the Art
Model LLM Frames PerceptionTest NExT-QA EgoSchema VNBench LongVideoBench Video-MME MLVU
Duration 23 sec 44 sec 180 sec 180 sec 8 sec - 60 min 1 - 60 min 3 - 120 min
@@
LLoVi [79] GPT3.5 - - 66.30 52.00 - - - -
VideoAgent [67] GPT4 - - 71.30 54.10 - - - -
VideoTree [69] GPT4 - - 75.60 61.10 - - - -
@@
LLaVA-NeXT (Video) [44] Vicuna-7B 64 46.13 67.66 41.66 - - 42.21 42.33
PLLaVA [75] Vicuna-7B 64 48.55 67.56 43.36 - - 42.13 44.61
BIMBA-LLaVA Vicuna-7B 64 52.61 72.35 52.31 - - 45.66 47.16
LLaMA-3.2 (Video) [52] LLaMA3.2-8B 64 53.16 73.72 55.21 - - 47.31 47.61
PLLaVA (LLaMA-3.2) [52] LLaMA3.2-8B 64 53.13 72.77 52.33 - - 46.56 46.56
BIMBA-LLaMA LLaMA3.2-8B 64 56.50 76.88 60.25 - - 50.11 50.60
@@
LLaMA-VID [41] Vicuna-7B 1fps 44.60 - 38.50 10.80 - - 33.20
Video-LLaVA [42] Vicuna-7B 8 44.30 62.60 38.40 - - 40.40 47.30
LLaVA-NeXT-Video [82] Vicuna-7B 32 48.80 - 43.90 20.10 43.50 46.50 -
PLLaVA [75] Vicuna-7B 16 - 68.17 45.16 - 40.20 44.25 47.31
VideoChat2 [39] Vicuna-7B 16 47.30 - 54.40 12.40 36.00 47.90 54.60
LongVA [81] Qwen2-7B 128 - 68.30 - 41.50 - 54.30 56.30
Video-LLaMA2 [10] Qwen2-7B 32 51.40 - 51.70 24.90 - 47.90 48.50
LLaVA-OneVision [34] Qwen2-7B 32 57.10 79.40 60.10 51.80 56.5 58.20 64.70
Kangaroo [47] LLaMA2-8B 64 - - 62.70 - - 56.00 61.00
Video-XL [62] Qwen2-7B 2048 - - - 61.60 49.50 55.50 64.90
LongVU [61] Qwen2-7B 1fps - - 67.60 - - 60.60 65.40
Qwen2-VL [66] Qwen2-7B 2fps - - - - 55.60 63.30 -
LLaVA-Video [83] Qwen2-7B 64 67.90 83.20 57.30 70.77 58.20 63.30 70.80
BIMBA-LLaVA Qwen2-7B 128 68.51 83.73 71.14 77.88 59.46 64.67 71.37
@@
Table 3: We compare BIMBA with state-of-the-art video MLLMs across seven diverse video question-answering benchmarks. BIMBA-LLaVA achieves the highest performance on all datasets when using the Qwen2-7B LLM backbone (third section). Since different MLLMs leverage varying LLM backbones and training data, we also conduct a fair comparison by evaluating our model against four baselines trained using the same 370K instruction-tuning dataset and using Vicuna-7B and LLaMA3.2-8B LLM decoders (second section). In this evaluation setting, we observe that our model consistently outperforms all baselines across all benchmarks.
@@
In this section, we compare our model against state-of-the-art video MLLMs on seven different video benchmarks, summarizing the results in Table 3. Since prior models rely on varying backbones and training datasets, to ensure a fair comparison, we first fine-tune four competitive models, as well as the two default variants of our BIMBA model on our 370K instruction-tuning dataset. These include PLLaVA [75], LLaVA-NeXT (Video) [44], LLaMA-3.2 [52] (Video) and PLLaVA (LLaMA-3.2), a variant of PLLaVA using the stronger LLaMA-3.2 [52] image MLLM. The results in the second section of the table show that our model achieves better results than all these baselines. In particular, we note that when using the same Vicuna-7B [11] model, BIMBA-LLaVA outperforms both PLLaVA and LLaVA-NeXT (Video) on all benchmarks, achieving an average accuracy of 55.23% compared to 49.42% for LLaVA-NeXT and 50.40% for PLLaVA. Likewise, our BIMBA-LLaMA variant surpasses LLaMA-3.2 (Video) by 3.09% on average and PLLaVA (LLaMA-3.2) by 4.21%. These results highlight the effectiveness of our selective-scan-based token compression mechanism, which can be seamlessly integrated into existing MLLMs to enhance performance across a wide range of video understanding tasks.
@@
Furthermore, since recent MLLMs [34, 47, 62, 61, 66, 83] have been trained with significantly larger instruction-tuning datasets and stronger LLM backbones, we also trained a variant of BIMBA-LLaVA using approximately 1.6 million video instruction samples and the Qwen2-7B LLM backbone. The results in the third section of Table 3 show that this BIMBA-LLaVA variant achieves state-of-the-art performance on multiple challenging video question-answering benchmarks, including NExT-QA [73], EgoSchema [51], VNBench [85], LongVideoBench [71], and Video-MME [15]. On the EgoSchema benchmark, our model surpasses the previous best method, LongVU [61], by 3.54%, demonstrating its superior ability to comprehend egocentric videos and handle questions that require long context understanding. Similarly, on VNBench, which focuses on needle-in-the-haystack questions, our approach outperforms LLaVA-Video [83] by 7.11%, highlighting its strong capability to extract key information from very long videos. Additionally, BIMBA-LLaVA (Qwen2-7B) substantially outperforms models based on proprietary LMMs. Specifically, it surpasses VideoTree [69] (GPT-4) by 7.7% on NExT-QA and 10.04% on EgoSchema. Finally, on benchmarks requiring long video comprehension, such as LongVideoBench [71], Video-MME [15], and MLVU [86] our model sets a new state-of-the-art, further demonstrating its effectiveness in processing and understanding hour-long videos.
@@ 7 Conclusion
Acknowledgements. We thank Yan-Bo Lin, Feng Cheng, Ce Zhang, Yue Yang, Yulu Pan, Han Yi, Triantafyllos Afouras, and Effrosyni Mavroudi for their helpful discussions. This work was supported by the Laboratory for Analytic Sciences via NC State University, ONR Award N00014-23-1-2356.
@@
005050505010010010010015015015015000101010102020202030303030SecondsAccuracy Improvement
(a) LLaVA Backbone on NeXT-QA.
0050505050100100100100150150150150005555101010101515151520202020SecondsAccuracy Improvement
(b) LLaMA Backbone on NeXT-QA.